Verify Rank-nullity theorem for the linear transformation T:R³----->R³ defined by T(x,y,z)=(x-y, 2y+z, x+y+z).

1
Expert's answer
2021-12-14T07:58:32-0500

Let us verify Rank-nullity theorem for the linear transformation "T:\R^3\to \R^3" defined by "T(x,y,z)=(x-y, 2y+z, x+y+z)."


For this let us find the kernal and the range of "T."


"Ker(T)=\{(x,y,z)\in\R:T(x,y,z)=(0,0,0)\}
\\=\{(x,y,z)\in\R:(x-y, 2y+z, x+y+z)=(0,0,0)\}
\\=\{(x,y,z)\in\R:x-y=0, 2y+z=0, x+y+z=0\}
\\=\{(x,y,z)\in\R:x=y, z=-2y, x+y+z=0\}
\\=\{(y,y,-2y)\in\R:y\in\R\}
\\=\{y(1,1,-2)\in\R:y\in\R\}."


It follows that "\dim(Ker(T))=1."


Let "(a,b,c)\in Range(T)." Then "T(x,y,z)=(a,b,c)" for some "(x,y,z)\in\R."

Therefore, "x-y=a, 2y+z=b, x+y+z=c."

It follows that "x-y=(x+y+z)-(2y+z)=c-b," and hence "a=c-b."

Then


"Range(T)=\{(a,b,c)\in\R: T(x,y,z)=(a,b,c)\}
\\=\{(c-b,b,c):b,c\in\R\}
\\=\{c(1,0,1)+b(-1,1,0):b,c\in\R\}."


Consequently, "\dim(range(T))=2."


Since "\dim(Ker(T))+\dim(range(T))=1+2=3=\dim(\R^3)," we conclude that the Rank-nullity theorem is indeed true.